If the market interest rate is 5 percent, what is the present discounted value of a financial instrument that pays you $50 per year, forever, starting next year, with the exception of year 17. I.e. it pays you $50 every year except for the payment 17 years from today, which is zero.  
Answer #1

ANSWER:

I = 5%

PV = Payment per year / interest rate - payment in 17th year not paid(p/f,i,n)

pv = 50 / 5% - 50(p/f,5%,17)

pv = 1,000 - 50 * 0.4363

pv = 1,000 - 21.815

pv = 978.185

so the present discounted value is $978.185